OpenStudy (anonymous):

express in a+bi form : 2-3i 5+i

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

\[\frac{2-3i}{5+1}\times \frac{5-i}{5-i}=\frac{(2-3i)(5-i)}{5^2+1^2}\] is a start

OpenStudy (anonymous):

the denominator is a real number, namely \(26\) and the numerator is what you get when you multiply out